OpenStudy (anonymous):

Find a function whose graph is a parabola with vertex (4,-23) and that passes through the point (5,-21). Please show work.

OpenStudy (dumbcow):

vertex form \[y = a(x-h)^{2} + k\] vertex: (h,k) h = 4 k = -23 \[y = a(x-4)^{2} - 23\] Now use the point (5,-21) to solve for a substitute 5 for x, -21 for y \[-21 = a(5-4)^{2} -23\] \[-21 = a - 23\] \[a = 2\] Then the function is \[y = 2(x-4)^{2} - 23\]